OpenStudy (anonymous):

Find the area of a rectangle whose dimension is: 48 square root of 6 feet by 50 square root of 3 feet. How?

OpenStudy (anonymous):

Area of a rectangle is just multiplying the side lengths. So it's just \(48\sqrt{6} \times 50\sqrt{3}\)

OpenStudy (anonymous):

so the answer would be 2400 square root of 18? is it right?

OpenStudy (anonymous):

Yes. You can simplify the \(\sqrt{18}\) part if you want to (or have to) but it's not wrong the way it is either.

OpenStudy (anonymous):

how can i simplify the square root of 18? i really don't know.

OpenStudy (anonymous):

\[\sqrt{18} = \sqrt{9 \times 2} = \sqrt{9} \times \sqrt{2} = 3\sqrt{2}\]

OpenStudy (anonymous):

By remembering that \(\sqrt{a} \times \sqrt{b} = \sqrt{a \times b}\) you can find factors that are square numbers and simplify square roots like that.

OpenStudy (anonymous):

what is the answer when the square root is already simplified? 'm really confused

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

i mean that, my answer was 2400 square root of 18, then you told me that the square root of 18 can be simplified. my question is what is the answer now if the square root of 18 is already simplified?

OpenStudy (anonymous):

It's just your old answer times the simplified version... \[2400\sqrt{18} = 2400 \times 3\sqrt{2} = 7200\sqrt{2}\]

OpenStudy (anonymous):

oh, now i know. Thanks for the great help. :)

OpenStudy (lukecrayonz):

Che, imagine 2400sqrt(18) remove sqrt(18) away. Now, you know that 18 can be broken down as things can multiply to it. 9*2=18 Now break 9 down. 3*3=9 So now 3*3*2=18. Because it is a square root, you can take out pairs. You have two 3's, so now: 3sqrt(2)=sqrt(18) 2400*3sqrt(2)=7200sqrt(2) as you can multiply the outside numbers

OpenStudy (anonymous):

oh, i already understand, thanks for the explanation.